trigonometry how to prove that: sin 2x = 2 sinx cosx

trigonometry
how to
prove that:
 sin 2x =  2 sinx cosx

Grade:11

1 Answers

Mehdi Alam
5 years ago
sin(x+y)=sinx.cosy+cosx.siny,if we put x=y, then, Sin(x+y)=sin(2x)&=sinx.cosy+cosx.siny=sinx.cosx+cosx.sinx =2sinx.cosx£, sin(2x)=2sinx.cosx